40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Jul 23 15:00:28 2017, Jamie, Omnistructure, VAC, strange sound around X arm vacuum pumps 
    Reply  Mon Jul 24 09:55:29 2017, Steve, Omnistructure, VAC, controller failer TP2controllerFails.png
       Reply  Tue Jul 25 08:48:57 2017, Steve, Update, VAC, RGA scan at d278 RGA278d.png
          Reply  Mon Aug 28 10:28:07 2017, Steve, Update, VAC, RGA scan 20 day after 17 Torr 20d_after_17_torr.png
             Reply  Tue Jan 30 10:22:20 2018, Steve, Update, VAC, RGA scan at day 175 RGApd80-b-468-175d.pngRgaOn20hrs.png
                Reply  Fri Mar 16 09:36:56 2018, Steve, Update, VAC, RGA scan at day 511,218d RgaScan511d.png08.png
                   Reply  Mon Mar 26 09:25:18 2018, Steve, Update, VAC, VM1 opened CC4VM1.png
                      Reply  Tue Mar 27 08:58:21 2018, Steve, Update, VAC, VM1 opened.......scan fine rga2018march27.png
          Reply  Tue Sep 19 11:05:26 2017, Steve, Update, VAC, RGA scan at day 334  RGAscan334d.png
             Reply  Fri Oct 27 10:14:11 2017, Steve, Update, VAC, RGA scan at day 372 RGA_scan_d372.png
                Reply  Mon Oct 30 10:09:41 2017, Steve, Update, VAC, TP2 failed tp2failed.pngifo_1.0E-5_Torrit.pngtp2failed2dago.png4days.png
                   Reply  Mon Nov 6 09:09:43 2017, Steve, Update, VAC, setting up new TP2 turbo 
       Reply  Tue Jul 25 14:04:06 2017, Steve, Update, VAC, turbo controller installed and we are running at vac normal TP2@75Krpm.png
Message ID: 13407     Entry time: Mon Oct 30 10:09:41 2017     In reply to: 13403     Reply to this: 13409
Author: Steve 
Type: Update 
Category: VAC 
Subject: TP2 failed 

 IFO pressure 1.2e-5 Torr at 9:30am

Quote:

Valve configuration: Vacuum normal

Note: Tp2 running at 75Krpm 0.25A 26C has a  load high pitch sound today. It's fore line pressure 78 mTorr. Room temp 20C

 

Atm. 1,  This was the vacuum condition  this morning.

               IFO P1  9.7 mTorr, V1 openV4 was in closed position , ~37 C warm Maglev at normal 560Hz rotation speed with foreline pressure 3.9 Torr because V4 closed 2 days ago when TP2 failed .....see Atm.3

               The error messege at TP2  controller was: fault overtemp.

I did the following to restored IFO pumping: stopped pumping of the annulose with TP3 and valves were configured so TP3 can be the forepump of the Maglev.

closed VM1 to protect the RGA,  close PSL shutter .....see Gautam  entry

aux fan on to cool down Maglev-TP1, room temp 20 C,

aux drypump turned on and opend to TP3 foreline to gain pumping speed,

closed PAN to isolate annulos pumping,

opened V7 to pump Maglev forline with TP3 running at 50Krpm, It took 10 minutes to reach P2  1mTorr from 3.9 Torr

aux drypump closed off at P2  1 mTorr, TP3 foreline pressure 362 mTorr.......see Atm.2

As we are running now:

IFO pressure 7e-6 Torr at Hornet cold cathode gauge at 15:50  We have no IFO CC1 logging now.  Annuloses are in 3-5 mTorr range are not pumped.

TP3 as foreline pump of TP1 at 50 Krpm, 0.24 A, 24 C, it's drypump forline pressure 324 mTorr

V4 valve cable is disconnected.

I need help with wiring up the logging of the Hornet cold cathode gauge.

 

 

 

Attachment 1: tp2failed.png  24 kB  Uploaded Mon Oct 30 11:10:49 2017  | Hide | Hide all
tp2failed.png
Attachment 2: ifo_1.0E-5_Torrit.png  24 kB  Uploaded Mon Oct 30 11:11:20 2017  | Hide | Hide all
ifo_1.0E-5_Torrit.png
Attachment 3: tp2failed2dago.png  33 kB  Uploaded Mon Oct 30 12:26:55 2017  | Hide | Hide all
tp2failed2dago.png
Attachment 4: 4days.png  45 kB  Uploaded Wed Nov 1 10:35:09 2017  | Hide | Hide all
4days.png
ELOG V3.1.3-